OpenStudy (anonymous):

Do the dollars earned form an arithmetic sequence? If so, determine the formula. Hours Worked 1 2 3 4 Dollars Earned 48.50 57.00 65.50 74.00

OpenStudy (anonymous):

An arithmetic sequence is a sequence where every term is dependent on the one before, and the difference between any two following terms are the same. Is this true in this case?

OpenStudy (anonymous):

Basically, the numbers 48.50, 57.00, 65.50, and 74.00 form a sequence. You have to see if there is a common difference. If the following is true, then it does form an arithmetic sequence. \[57 - 48.5 = 65.5 - 57 = 74 - 65.5\]Is this statement true when simplified?